OpenStudy (anonymous):

How many milliliters equal 1942 cm3?

OpenStudy (anonymous):

a.1.942 b.194.2 c.1942 d.19,420

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

If \[\Large 1 cm^3 = 1 mm \] then \[\Large 1942 \times 1 cm^3 = 1942 \times 1 mm\]

OpenStudy (anonymous):

yeas

OpenStudy (agent0smith):

Oops, those should be mL not mm.

OpenStudy (agent0smith):

So...can you find the answer from that?

OpenStudy (anonymous):

prob not

OpenStudy (agent0smith):

\[\Large 1942 \times 1 cm^3 = 1942 \times 1 mL\] what is 1942 times 1?

OpenStudy (anonymous):

1942

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

is it 1942?

OpenStudy (anonymous):

medal wil l be given!!!!!!!!!!!!!!!!!!11

OpenStudy (anonymous):

plzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzzz

OpenStudy (anonymous):

I got a different answer than 1942

OpenStudy (anonymous):

wat did u get

OpenStudy (anonymous):

oops never mind that is what I got Sorry : )

OpenStudy (anonymous):

Sorry I did it wrong first time! :)

OpenStudy (agent0smith):

\[\Large 1942 \times 1 cm^3 = 1942 \times 1 mL\] then \[\Large 1942 cm^3 = 1942 mL\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohh lol, so that it?

OpenStudy (anonymous):

yeah you were right it's 1942ml

OpenStudy (anonymous):

thanks

OpenStudy (agent0smith):

@vvvv do you understand it? How many milliliters equal 500 cm^3?

OpenStudy (anonymous):

500

OpenStudy (agent0smith):

correct :) \[\Large 500 \times 1 cm^3 = 500 \times 1 mL\] then \[\Large 500 cm^3 = 500 mL\]

OpenStudy (anonymous):

yae

OpenStudy (anonymous):

thanks